# Energy shifts and broadening of excitonic resonances in electrostatically-doped semiconductors

### v1

@@ IV.2 Energy blueshift of neutral excitons through evacuation of distinguishable resident carriers
We demonstrate how this physics manifest through the absorption resonances of neutral excitons in WSe2 monolayers. The choice to focus on absorption resonances in this monolayer is intentional because we can then readily decipher the roles of distinguishability and optimality in all three possible configurations of the exciton under various conditions. On the other hand, the recombination and energy relaxation of neutral excitons involve exchange scattering processes that complicate the analysis of emission resonances [18]. Exchange scattering is also relevant when dealing with absorption resonances of excited-state neutral excitons, and we leave the analysis of this case to Appendix E.
@@
Figure (12)(a) shows the energy shift and FWHM of the neutral exciton absorption resonance as a function of charge density in an hBN-encapsulated WSe2 monolayer at zero magnetic field. These results were extracted from low-temperature reflectance spectra, where details of the sample fabrication and experimental setup can be found in the supplementary information of Ref. [60]. When the holes density increases from zero to 1012 cm-2, the energy blueshift is 7.5 meV and the FWHM increases from 2 to 6 meV. Half of the resident holes are distinguishable and the other half are indistinguishable, as shown by the left diagram of Fig. (12)(b). When the monolayer is electron-doped, the corresponding energy blueshift is 10 meV and the FWHM increases from 2 to 7 meV. The resident electrons are 100% distinguishable in this case, as shown by the right diagram of Fig. (12)(b). Assuming that distinguishability of the resident carriers is the key point regardless of whether these are electrons or holes (as long as their effective masses are comparable), we can use the results of Fig. (12)(a) to infer the following. Increasing the density of distinguishable (indistinguishable) resident carriers by 1011 cm-2 roughly adds 1 meV (0.5 meV) to the blueshift and 0.5 meV (0.3 meV) to the broadening of the neutral exciton resonance. That is, the effect from distinguishable resident carrier is about as twice as strong.
@@
The stronger effect caused by the presence of distinguishable resident carriers is corroborated by helicity-resolved measurements of the energy shift and FWHM from the same device. Figures 12(c) and (d) show the respective results when $B=9\,$T, where diagrams of the corresponding absorption resonances are shown in Figs. (12)(e). We can confirm that the holes are fully valley polarized through the steady energy redshift and constant FWHM of the distinct resonance of the positive trion, labeled by $X^{+}(\sigma_{-})$. Focusing on the neutral exciton diagrams in Fig. 12(e), the valley-polarized resident holes are 100% distinguishable when dealing with the distinct resonance of the suboptimal exciton, labeled by $X^{0}(\sigma_{-})$, and 100% indistinguishable when dealing with the indistinct resonance of the optimal exciton, labeled by $X^{0}(\sigma_{+})$. Comparing the energy blueshifts of $X^{0}(\sigma_{\pm})$ in Fig. (12)(c), we again notice that distinguishable resident carriers induce as twice as strong energy blueshift. The FWHM results in Fig. (12)(d) show that the broadening rate of $X^{0}(\sigma_{+})$ is strongly suppressed compared with that of $X^{0}(\sigma_{-})$. Given that there are no distinguishable resident carriers with which an optimal exciton can form a trion, the broadening of $X^{0}(\sigma_{+})$ is not limited by the same ultrafast process that converts suboptimal excitons to optimal trions (i.e., $X^{0}(\sigma_{-})\,\rightarrow X^{+}(\sigma_{-})$).
Finally, to be consistent with the analysis in Sec. IV.1, we also explain how the energy blueshift of $X^{0}$ emerges in case that the resident carriers are localized. When the density of distinguishable resident carriers increases, the energy of the exciton has to increase as well (i.e., to blueshift) for its state to remain extended, or else the exciton binds to a localized electron and becomes a localized trion [62, 61]. When the density of indistinguishable resident carriers increases, the energy of the exciton has to increase, albeit at a lesser blueshift rate, because of the smaller region left in the sample for the exciton to form [61].
V Compressibility and exciton resonances
@@
Similarly, exciton formation is inhibited in a monolayer subjected to LL quantization at integer filling. Without loss of generality, we explain this physics in a hole-doped monolayer (e.g., Figs. 13(a) and (c)). The hole of the exciton is not subjected to LL quantized motion because the exciton is charge neutral. At integer filling conditions, the holes in the LLs cannot make room for the indistinguishable hole of the exciton (with similar spin and valley). Having a partially filled LL alleviates this constraint because holes from this LL can hop between cyclotron orbit centers. That is, indistinguishable resident holes avoid the exciton by elastic scattering away from the photoexcited hole.
Finally, we notice that the energy blueshift of $X^{0}$ along the dotted black lines in Figs. 13(a),  13(c),  14(c) and  14(d) are all related to the density of compressible resident carriers. This dependence is implicit in the case of a partially filled LL, denoted by the fraction term in Eq. (5), where the density of resident carriers in a LL is proportionate to the magnetic field. This dependence in explicit in the moiré heterobilayer system, evidenced from the fraction term in Eq. (6), which denotes the relative density of itinerant resident carriers between adjacent incompressible states. Deriving the direct relation between compressibility and energy shift, through which one can better quantify the phenomenological $C$ terms in Eqs. (5)-(6), will be studied elsewhere.
VI conclusion
@@
We have self-consistently analyzed the various trends of energy shifts and broadenings of neutral excitons in their ground and excited states. The stronger energy blueshift and broadening of suboptimal excitons originate from a shakeup process that evacuates distinguishable resident carriers from the immediate vicinity of the exciton. The neutral exciton shows resilience in the absence of distinguishable resident carriers (i.e., smaller energy shift, weaker broadening, and stronger amplitude), rendering the exciton an optimal complex, and especially in moiré heterobilayer systems wherein its resonance can also become distinct. Finally, we have elucidated the intimate connection of the compressibility of the resident carrier gas to the amplitude and energy shift of the neutral exciton resonance.
@@ Acknowledgements.
Appendix A Hexcitons and oxcitons
@@
Much of the way we perceive excitons in semiconductors is attributed to the seminal work of Elliott, who tailored the hydrogen model to calculate exciton states [63]. Employing the semiconductor’s dielectric constant and effective masses of its CB electron and VB hole, the two-body model successfully predicts the energy spectrum of charge-neutral excitons when using appropriate Coulomb potential models [65, 64, 66]. Similar analogy can be drawn between a trion and hydrogen ion, H-, where the trion complex is modeled as a three-body problem [67, 68, 69, 70, 71]. Solutions of this problem yield binding energies which agree with the measured energy difference between the trion and exciton resonances in the spectra of TMD monolayers [72, 73, 74, 75, 76, 77, 78].
@@
If we wish to generalize this idea further to consider larger excitonic complexes, the following subtlety should be borne in mind. The hydrogen ion $H^{-N}$, where $N$ is the number of extra electrons, is unstable when $N>1$. One proton cannot hold more than two electrons, and therefore $H^{-N}$ dissociates to $H^{-}$ and $N-1$ electrons that end up far away from the ion and from each other. On the other hand, the stability of an excitonic complex is not a problem when dealing with an electrostatically-doped semiconductor that has $N$ Fermi seas. Here, the photoexcited pair acts as a charge-neutral hydrogen atom, and it bind to $N$ electrons with different spin-valley configurations which are near each other before excitation. To explain this point, let us consider a semiconductor whose CB has $N=2$ populated valleys and the photoexcited pair is distinguishable, as shown in Fig. 2(d). The average distance between nearby electrons, one from the valley at $K$ and the other from $-K$, is about $k_{F}^{-1}$ where $k_{F}$ is the Fermi wavenumber. When the electron density is large enough, the hexciton formation means that the photoexcited pair binds to these electrons which are already near each other at equilibrium. The same principle holds true when $N=3$ (oxciton) [24] or in multivalley semiconductors with $N\geq 4$.
@@
Appendix B LL quantization of hexciton and oxciton resonances in a strong magnetic field
@@
This behavior supports the observed LL branches in Fig. 15(b) and the observation that the resonance $H_{i}$ continues to redshift despite being indistinct (Sec. IV.1). Namely, the photoexcitation does not require a shakeup process to accommodate the Pauli exclusion principle if the optically-active electron is promoted to an empty LL level and its quantized motion resembles that of a free electron.
@@
Appendix C Alternative explanation to the universal energy redshift

### v2

@@ V Energy shift of neutral excitons
We demonstrate and quantify the energy blueshift of the absorption resonances of neutral excitons in WSe2 monolayers. The choice to focus on absorption resonances in this monolayer is intentional because we can then readily decipher the roles of distinguishability and optimality in three possible configurations of the exciton under various conditions. On the other hand, the recombination and energy relaxation of neutral excitons involve exchange scattering processes that complicate the analysis of emission resonances [19]. Exchange scattering is also relevant when dealing with absorption resonances of excited-state neutral excitons, and we leave the analysis of this case to Appendix I.
@@
Figure 12(a) shows the energy shift and FWHM of the neutral exciton absorption resonance as a function of charge density in an hBN-encapsulated WSe2 monolayer at zero magnetic field. These results were extracted from low-temperature reflectance spectra, where details of the sample fabrication and experimental setup can be found in the supplementary information of Ref. [88]. When the hole density increases from zero to 1012 cm-2, the energy blueshift is 7.5 meV and the FWHM increases from 2 to 6 meV. Half of the resident holes are distinguishable and the other half are indistinguishable, as shown by the left diagram of Fig. 12(b). When the monolayer is electron-doped, the corresponding energy blueshift is 10 meV and the FWHM increases from 2 to 7 meV. The resident electrons are 100% distinguishable in this case, as shown by the right diagram of Fig. 12(b). Assuming that distinguishability of the resident carriers is the key point regardless of whether these are electrons or holes (as long as their effective masses are comparable), we can use the results of Fig. 12(a) to infer the following. Increasing the density of distinguishable (indistinguishable) resident carriers by 1011 cm-2 roughly adds 1 meV (0.5 meV) to the blueshift and 0.5 meV (0.3 meV) to the broadening of the neutral exciton resonance. That is, the effect from distinguishable resident carrier is about as twice as strong.
@@
The stronger effect caused by the presence of distinguishable resident carriers is corroborated by helicity-resolved measurements of the energy shift and FWHM from the same device. Figures 12(c) and (d) show the respective results when $B=9\,$T, where diagrams of the corresponding absorption resonances are shown in Fig. 12(e). We can confirm that the holes are fully valley polarized through the steady energy redshift and constant FWHM of the distinct resonance of the positive trion, labeled by $X^{+}(\sigma_{-})$. Focusing on the neutral exciton diagrams in Fig. 12(e), the valley-polarized resident holes are 100% distinguishable when dealing with the distinct resonance of the suboptimal exciton, labeled by $X^{0}(\sigma_{-})$, and 100% indistinguishable when dealing with the indistinct resonance of the optimal exciton, labeled by $X^{0}(\sigma_{+})$. Comparing the energy blueshifts of $X^{0}(\sigma_{\pm})$ in Fig. 12(c), we again notice that distinguishable resident carriers induce as twice as strong energy blueshift. The FWHM results in Fig. 12(d) show that the broadening rate of $X^{0}(\sigma_{+})$ is strongly suppressed compared with that of $X^{0}(\sigma_{-})$. Given that there are no distinguishable resident carriers with which an optimal exciton can form a trion, the broadening of $X^{0}(\sigma_{+})$ is not limited by the same ultrafast process that converts suboptimal excitons to optimal trions (i.e., $X^{0}(\sigma_{-})\,\rightarrow X^{+}(\sigma_{-})$). These results are consistent with previous findings [42, 7].
V.1 Indistinguishable resident carriers
[…]
The dominant role of the exchange interaction between an exciton and resident carriers in semiconductors can facilitate such optical process, wherein the electron (hole) component of the exciton is exchanged with a resident electron (hole) after their interaction [49, 19]. We elaborate on this mechanism in Appendix H, where we also show that the exchange scattering of the exciton with resident carriers is far stronger than the exciton-phonon interaction or than the direct interaction, wherein the components of the exciton are kept the same before and after scattering. Figure 13(c) shows the calculated absorption rate when the exciton creation is accompanied by exciton-hole exchange interaction (see Appendix H for details). The energy blueshift and broadening match the trends seen in experiment, wherein absorption at a larger energy involves larger momentum transfer during the exciton-hole exchange interaction.
Figure 13: (a) Absorption diagram of a suboptimal neutral exciton with indistinct resonance. (b) The optical process is facilitated by the exciton-electron exchange interaction (see text and Appendix  H). (c) Rates of exchange-assisted exciton absorption in hBN-encapsulated WSe2 monolayer, calculated in the presence of indistinguishable resident holes at four densities (solid lines). The calculated rates are normalized with respect to the direct exciton absorption rate at zero charge density (dotted line), whose resonance is the reference energy.
V.2 Distinguishable resident carriers
Band filling by distinguishable resident carriers does not Pauli block the optical transition, and their presence does not Pauli exclude the photoexcited e-h pair. Yet, we have seen that the energy blueshift of suboptimal excitons is strongest when the semiconductor only hosts distinguishable resident carriers (Fig. 12). The presence of these carriers means that there is an optimal complex at a lower energy than the suboptimal exciton. In connection with the scattering phenomenon we have discussed before, the suboptimal exciton refers to a case that the photoexcited e-h pair scatters off a distinguishable resident carrier rather than binds to it. To study the energy blueshift effect of the suboptimal exciton and the continuous transfer of its oscillator strength to the optimal complex at lower energy, we will render the relation between the scattering phase shift and excitonic density of states (DOS). This relation allows one to understand the relatively strong energy blueshift of the suboptimal complex by quantifying how the DOS function of the exciton-electron system evolves when the density of distinguishable resident carriers increases.
Figure 14: (a) Excitonic density of states (DOS) function in a two-dimensional semiconductor with and without distinguishable resident carriers. $\varepsilon_{\text{1s}}$ and $\varepsilon_{\text{b,T}}$ are the resonance energy of the suboptimal exciton and binding energy of the optimal trion, respectively. (b) Calculated excitonic DOS function in electron-doped WSe2 when the electrons form a square lattice and their charge density is $5\times 10^{11}$ cm-2 (solid green line) and $2\times 10^{12}$ cm-2 (dashed-dotted black line). The latter is also calculated when the distinguishable electrons are randomly distributed in the sample (dashed magenta line). The dotted cyan line shows the excitonic DOS function without resident electrons, where $\varepsilon_{\text{1s}}$ is the zero reference energy. (c) Calculated density dependence of the energy blueshift of excitons when the distinguishable electrons form a square Wigner lattice (black solid line), indistinguishable and form a square Wigner lattice (red solid line), or indistinguishable and randomly distributed (red dashed line). Taken from Ref. [93].
Let us consider an effective two-body system comprising a distinguishable resident (itinerant) electron and a point-like exciton. The relative motion between the exciton and electron is governed by a dipole-charge interaction, which if approximated by a central potential, the change to the DOS function for each partial wave can be written as [89],
$\Delta D_{l}(\varepsilon)=\begin{cases}\sum_{j}\delta(\varepsilon-\varepsilon_{j})\,\,,&\varepsilon<0\\ \\ {{\frac{1}{\pi}\frac{\delta_{l}}{d\varepsilon}}}\,\,,&\varepsilon>0\,.\end{cases}$ (6)
The sum in the first line runs over bound states (i.e., trion states), and the second line corresponds to continuum states (the exciton and resident carrier are not bound to each other). The phase shift $\delta_{l}$ is introduced by the scattering between the exciton and the electron in this case. Focusing on the optically active $s$ states ($l=0$), the Levinson’s theorem implies that [89, 90, 91, 92]
[…]
This is essentially a zero sum rule which states that adding to the DOS in one spectral region mandates taking away from the DOS in a different spectral region. Relevant to the energy blueshift of the suboptimal exciton resonance, forming a bound trion state at $\varepsilon<0$ implies depleting the low-energy DOS region of the exciton at $\varepsilon>0$.
Figure 14(a) shows characteristic schemes of the DOS function with (solid black line) and without (dotted red line) distinguishable resident carriers. The DOS function units are $M_{x}/\pi\hbar^{2}$, where $M_{x}$ is the translational mass of the exciton in a two-dimensional semiconductor. The exciton resonance energy is $\varepsilon_{\text{1s}}$ and the binding energy of the trion is $\varepsilon_{\text{b,T}}$. The transfer of spectral weight from the low energy part of the step function (slightly above $\varepsilon_{\text{1s}}$) to the trion resonance depends on the details of the scattering phase shift. This transfer gives rise to the energy blueshift of the exciton and to the transfer of oscillator strength from the exciton to the trion [93]. Such transfer does not arise when the semiconductor is only populated by indistinguishable resident carriers, in which case the exciton is optimal and a trion state cannot be created.
The energy blueshift effect due to the transfer of DOS weight from the exciton to the trion spectral regions is a general effect that does not depend on the itinerancy of resident carriers. In case of localized and distinguishable resident carriers, we have used the recursion method to calculate the excitonic DOS function at various charge densities and order configurations. Technical details of these calculations can be found in Ref. [93], and here we present the final results. Figure 14(b) shows the excitonic DOS function of electron-doped WSe2 when the electrons are distinguishable and they form a square lattice. Their charge densities are $5\times 10^{11}$ cm-2 (solid green line) and $2\times 10^{12}$ cm-2 (dashed-dotted black line). The dotted cyan line shows the excitonic DOS function without resident electrons, where $\varepsilon_{\text{1s}}$ is the zero reference energy. The oscillatory behavior of the excitonic DOS function at positive energies is a result of the excitonic band structure when the electrons form a Wigner square lattice [94]. This oscillatory behavior is suppressed when the resident carriers are randomly distributed, as shown by the dashed magenta line in Fig. 14(b) for $2\times 10^{12}$ cm-2. The trion energy is broadened and its resonance amplitude weakens when the resident carriers are randomly distributed, but its resonance energy does not shift because the trion size is smaller than the average distance between resident carriers at these densities.
The solid black line in Fig. 14(c) shows the calculated energy blueshift of the exciton resonance as a function of the density of distinguishable resident electrons. To compare this blueshift with the case of localized indistinguishable resident electrons, the sign of the interaction between the electron and point-like dipole is switched [93], such that the dipole-charge potential is repulsive. The solid and dashed red lines in Fig. 14(c) show the energy blueshift of the exciton resonance as a function of the density of indistinguishable resident electrons when they form a Wigner square lattice or randomly distributed in the monolayer, respectively. Similar to the experimental result in Fig. 12, distinguishable resident electrons induce stronger energy blueshift, stemming from the presence of an optimal trion state at a lower energy.
VI Compressibility and exciton resonances
@@
Similarly, exciton formation is inhibited in a monolayer subjected to LL quantization at integer filling. Without loss of generality, we explain this physics in a hole-doped monolayer (e.g., Figs. 15(a) and (c)). The hole of the exciton is not subjected to LL quantized motion because the exciton is charge neutral [95, 96, 70]. At integer filling conditions, the holes in the LLs cannot make room for the indistinguishable hole of the exciton (with similar spin and valley). Having a partially filled LL alleviates this constraint because holes from this LL can hop between cyclotron orbit centers. That is, indistinguishable resident holes avoid the exciton by elastic scattering away from the photoexcited hole.
Finally, we notice that the energy blueshift of $X^{0}$ along the dotted black lines in Figs. 15(a),  15(c),  16(c) and  16(d) are all related to the density of compressible resident carriers. This dependence is implicit in the case of a partially filled LL, denoted by the fraction term in Eq. (8), where the density of resident carriers in a LL is proportionate to the magnetic field. This dependence in explicit in the moiré heterobilayer system, evidenced from the fraction term in Eq. (9), which denotes the relative density of itinerant resident carriers between adjacent incompressible states. Deriving the direct relation between compressibility and energy shift, through which one can better quantify the phenomenological $C$ terms in Eqs. (8)-(9), will be studied elsewhere.
VII conclusion
@@
We have self-consistently analyzed the various trends of energy shifts and broadenings of neutral-exciton absorption resonances. The stronger energy blueshift of excitons in the presence of distinguishable resident carriers is caused by the transfer of density-of-states from the suboptimal complex (exciton) to the optimal one (trion). The enhanced broadening in this case is attributed to the ultrafast time it takes a suboptimal exciton to bind to a distinguishable resident carrier. The neutral exciton shows resilience in the absence of distinguishable resident carriers (i.e., smaller energy shift, weaker broadening, and stronger amplitude), rendering the exciton an optimal complex, and especially in moiré heterobilayer systems wherein its resonance can also become distinct. Finally, we have elucidated the intimate connection of the compressibility of the resident carrier gas to the amplitude and energy shift of the neutral exciton resonance.
@@ Acknowledgements.
Appendix A Caveats
Quantum numbers that determine the distinguishability of an electron or hole in a semiconductor are its energy band, spin and wavevector ($n,s,\mathbf{k}$). To that one should add the monolayer index in case of heterobilayers. When dealing with excitonic complexes, we use the valley index of the electron (or hole) instead of its wavevector. The reason is that a particle can bind to a complex by moving with respect to other particles in the complex, and this internal motion requires a finite bandwidth of wavevectors (low energy states) in the valley at which the particle resides.
In general, the particles of an excitonic complex have to be mutually distinguishable, and they include resident electrons (holes) at the edge of the CB (VB) with different valley-spin configurations. Yet, there are exceptions to this rule. Biexcitons with two indistinguishable particles and two distinguishable ones can be stabilized only if the effective mass of the indistinguishable particles is at least five times larger than the effective mass of the distinguishable particles [47]. The relative motion of the lighter-mass distinguishable particles is much faster than that of the spatially non-overlapping indistinguishable particles, thereby suppressing the exchange effects between the heavy particles and keeping such biexciton bound.
The case of a bound trion that has two electrons (or holes) with the same spin and valley is more subtle. At zero or weak magnetic fields, such trion state can only emerge if the trion has nonzero angular momentum [48]. On the other hand, the optically active trions have zero angular momentum ($s$-states), and they cannot be stabilized with two particles that share the same spin and valley regardless of how heavy their mass may be. Such particles cannot be in contact with the opposite-charge particle simultaneously because of the Pauli exclusion principle, rendering the formation of a nodeless $s$-state impossible [49]. An exception to this rule can happen by applying a strong magnetic field, such that the separation between LLs (cyclotron energy) exceeds the trion binding energy. This condition can be met in semiconductor quantum wells, where the binding energy of the trion is less than 5 meV [50, 97, 51, 98, 99, 52, 100, 72, 69, 70, 71]. A trion state can then be stabilized even if it includes two particles with the same band-spin-valley configuration, because the LLs behave as additional distinguishability knob [69]. That is, the Pauli principle no longer provides a strong constraint as at zero field when the degeneracy of the LL is large enough [70].
Appendix B Hexcitons and oxcitons
@@
Much of the way we perceive excitons in semiconductors is attributed to the seminal work of Elliott, who tailored the hydrogen model to calculate exciton states [101]. Employing the semiconductor’s dielectric constant and effective masses of its CB electron and VB hole, the two-body model successfully predicts the energy spectrum of charge-neutral excitons when using appropriate Coulomb potential models [103, 102, 104]. Similar analogy can be drawn between a trion and hydrogen ion, H-, where the trion complex is modeled as a three-body problem [105, 106]. Solutions of this problem yield binding energies which agree with the measured energy difference between the trion and exciton resonances in the spectra of TMD monolayers [107, 108, 109, 110, 111, 112].
@@
If we wish to generalize this idea further, the following subtlety should be borne in mind when dealing with larger excitonic complexes. The hydrogen ion $H^{-N}$, where $N$ is the number of extra electrons, is unstable when $N>1$. One proton cannot hold more than two electrons, and therefore $H^{-N}$ dissociates to $H^{-}$ and $N-1$ electrons that end up far away from the ion and from each other. On the other hand, the stability of an excitonic complex is not a problem when dealing with an electrostatically-doped semiconductor that has $N$ Fermi seas. Here, the photoexcited pair acts as a charge-neutral hydrogen atom, and it binds to $N$ electrons with different spin-valley configurations which are near each other before excitation. To explain this point, let us consider a semiconductor whose CB has $N=2$ populated valleys and the photoexcited pair is distinguishable, as shown in Fig. 2(d). The average distance between nearby electrons, one from the valley at $K$ and the other from $-K$, is about $k_{F}^{-1}$ where $k_{F}$ is the Fermi wavenumber. When the electron density is large enough, the hexciton formation means that the photoexcited pair binds to these electrons which are already near each other at equilibrium. The same principle holds true when $N=3$ (oxciton) [25] or in multivalley semiconductors with $N\geq 4$.
@@
Appendix C LL quantization of hexciton and oxciton resonances in a strong magnetic field
@@
This behavior supports the observed LL branches in Fig. 17(b) and the observation that the indistinct resonance $H_{i}$ continues to redshift despite being indistinct (Sec. IV). Namely, the photoexcitation does not require a shakeup process to accommodate the Pauli exclusion principle if the optically-active electron is promoted to an empty LL level and its quantized motion resembles that of a free electron.
@@
Appendix D Energy shifts of optical gaps
The energy shift of a resonance, which one measures in experiment, reflects a change in the optical gap of the e-h pair that undergoes optical transition. When an excitonic complex is formed, it inherits the renormalized self energies of the resident carriers to which the e-h pair binds. Since these renormalized energies are essentially the same before and after the absorption, they cannot be measured through the absorbed photon energy.
To support this understanding, we focus on the Zeeman energy shift of the distinct resonance $X^{+}$ of the optimal positive trion in Figs. 15(b) and (d). This energy shift can be evaluated very well from $(g_{v}-g_{c})\mu_{B}B$, where $g_{c}=3.84$ and $g_{v}=6.1$ are the single-particle $g$ factors of the valleys that take part in the optical transition [12]. The exchange-induced renormalization energy of the hole to which the pair binds is not probed by the photon energy, as one can realize by the fact that the same slope is measured in Figs. 15(b) and (d), albeit the hole density is ramped up from $1.7\times 10^{12}$ in Fig. 15(b) to $4.6\times 10^{12}$ in Fig. 15(d). That is, the measured photon energy only reflects the rigid Zeeman shifts at the edges of the CB and VB valleys that take part in the optical transition.
This understanding can also be sharpened by contrasting this result with studies that report ‘enhanced’ $g$-factor values when the resident carriers are fully polarized [7, 57, 113]. For example, the ‘enhanced’ $g$-factor of holes in a fully valley-polarized WSe2 monolayer can be realized from the behavior of $X^{+}$ in Fig. 4(c). At $B=20\,$T, the holes in this device remain fully valley-polarized up to a density of $n_{h}=2.5\times 10^{12}$ cm-2 (corresponding to $V_{g}=-1.7$ V). This result means that the Zeeman energy difference between the ±$\pm K$ valleys, $E_{Z}=2g_{v}\mu_{B}B$, exceeds the Fermi energy $E_{F}=2\pi\hbar^{2}n_{h}/m_{h}$, where the holes populate one valley and their effective mass is $m_{h}=0.36m_{0}$ [55]. While the condition $E_{Z}\gtrsim E_{F}$ yields $g_{v}\gtrsim 14$, one measures $g_{v}\simeq 6$ from the energy shift of a trion resonance as a function of magnetic field [12]. The latter reflects a single-particle $g$-factor that one can also extract from first-principle calculations [114]. The conclusion is that the exchange interaction, which is responsible for the ‘enhanced’ $g$-factor through the self-energy of a resident hole [115], is not observed through the optical transition of an excitonic complex. This supports the understanding that the renormalized self-energies of resident carriers to which the e-h pair binds are essentially the same before and after the absorption, and therefore, they cannot be measured through the absorbed photon energy.
Appendix E Alternative explanation to the universal energy redshift